# Advancing General Multimodal Capability of Vision-language Models with Pyramid-descent Visual Position Encoding

### v1

@@ 1 Introduction
Large Language Models (LLMs) Touvron et al. (2023); Brown et al. (2020) demonstrate significant universal capabilities that contribute to the pursuit of general artificial intelligence. However, language constitutes only one aspect of communication. Visual information plays a crucial role in augmenting and enhancing our understanding of the world. Consequently, there is a growing interest in the development of Vision-language Models (VLMs) Chen et al. (2024c); Peng et al. (2023); Wang et al. (2024); Bai et al. (2023) that can simultaneously process and integrate various input modalities. To effectively leverage the powerful contextual understanding capabilities of LLMs, VLMs project visual information to the same dimensionality as textual embeddings through specific projection layers Chen et al. (2023); Li et al. (2023b); Zhou et al. (2024), which are then directly embedded into the text sequence to form the input for the foundation LLMs, enabling end-to-end cross-modal alignment and instruction-following learning using next-token prediction. Therefore, the image tokens are position-embedded together with the entire sequence.
Despite their commendable progress, the typical processing of visual information does not align with the distribution patterns of visual elements. Since visual information is composed of fixed-sized patches obtained through raster scanning, patches located closer to the bottom right corner of the image are positioned nearer to the instruction tokens within the sequence. Due to the long-term decay from Rotary Position Embedding (RoPE) Su et al. (2024), visual tokens closer to the instruction tokens will be more likely to receive higher attention weights, and vice versa. This is counterintuitive, as the importance of visual information is not defined by the order of raster-scanning. Xing et al. (2024) observe a similar phenomenon by visualizing the attention information flow from instruction tokens to visual tokens in the first layer of the Decoder. Consequently, they propose Concentric Causal Attention (CCA), which starts assigning the position indexes of images from the peripheral and ends in the center, to alleviate the limitations of long-term decay in RoPE and improve causal attention following 2D spatial locality of images. Although CCA is both intuitive and effective, its applicability is constrained by the assumption that all significant elements related to the instructions are situated at the center of the image. This assumption inherently results in a loss of detail, limiting its effectiveness in capturing comprehensive information.
@@ 2.1 Vision-language Model
Recent advancements in Vision-language Models (VLMs) have demonstrated impressive performance in processing multi-format information Huang et al. (2023); Achiam et al. (2023). VLMs are typically built upon existing Large Language Models (LLMs) and incorporate visual information as input tokens by utilizing an additional vision encoder (e.g., CLIP) and a bridging connector (e.g., MLP). For instance, LLaVA Liu et al. (2024a) employs an MLP to project visual tokens and aligns the feature dimensions with word embeddings, while BLIP-2 Li et al. (2023b) utilizes a set of learnable query tokens to extract information in a query-based manner. Building upon these foundational works, MM1 McKinzie et al. (2025) has further investigated the significance of the number of visual tokens and image resolution, identifying them as the most critical factors, while finding that the type of connector has minimal impact. By effectively connecting visual and textual modalities, VLMs significantly enhance human-AI interaction and exhibit remarkable capabilities in understanding and generating multimodal content Chen et al. (2024b); Peng et al. (2023); Chen et al. (2023); Wang et al. (2024); Hu et al. (2024); Xie et al. (2024).
@@ 2.2 Position Encoding for Transformers
Since transformer-based models contain no recurrence Hochreiter (1997) and convolution Islam et al. (2020) structure, additional information about the relative or absolute position of the tokens in the input sequence is required. Therefore, the community has witnessed the development of various position encoding methods, e.g. sinusoidal Vaswani (2017), learnable Dosovitskiy (2020), relative He et al. (2020); Shaw et al. (2018), and conditional Chu et al. (2021) position encoding. Among these studies, RoPE Su et al. (2024) are introduced to encode both absolute and relative positional information, showing superiority in pre-training LLMs Touvron et al. (2023); Achiam et al. (2023). The success of LLMs has led to the continued adoption of the effective RoPE scheme in VLMs for the unified encoding of positional information across sequences that incorporate multimodal features. However, it is important to note that visual information does not conform to the same one-dimensional sampling paradigm as language. Simple raster scanning is insufficient for modeling the spatial correlations among different patches. Consequently, numerous recent studies Chu et al. (2024); Xing et al. (2024); Lu et al. (2024) have sought to explore improved solutions that extend RoPE to visual tasks. In this paper, we investigate a novel multi-granularity position assignment strategy aimed at enhancing the VLM’s comprehension of visual information and improving the alignment between modalities.
@@ 3.1 Preliminaries
As indicated in Table 1, All-one PE performs weaker than the baselines and Concentric PE in perception but keeps competitive in coarse-grained perception tasks on different sizes of models. This suggests that, on the one hand, even when assigning the same positional weight to all image tokens, the VLM still possesses certain perception capabilities and performs better than raster-scan and concentric in coarse-grained perception. This is more pronounced on LLaVA-1.5-13B because larger models have stronger sequence modeling and feature capturing capabilities, which correspondingly bridge the gap in fine-grained abilities between All-one PE and other methods. On the other hand, the results from the baselines indicate that the long-term decay property of RoPE endows the image tokens with distinct weights, allowing the model to distinguish fine-grained tokens better and capture visual details.
@@ 3.2 Pyramid-descent Visual Position Encoding
Similar challenges were also present in the early development of Vision Transformer (ViT) Dosovitskiy (2020). Due to the columnar structure of ViT, which uses coarse image patches as input, it is difficult to apply it directly to pixel-level dense predictions such as object detection and segmentation. This difficulty arises because its output feature map is single-scale and low-resolution. To address these issues, Wang et al. (2021) proposed the Pyramid Vision Transformer (PVT). They utilize fine-grained image patches as input to learning high-resolution representations and introduce a progressive shrinking pyramid to reduce the sequence length of the Transformer as the network deepens, significantly lowering the computational cost. Moreover, compared to CNNs, PVT consistently produces a global receptive field, ensuring a holistic perception of visual elements and benefiting its performance in detection and segmentation tasks.
@@
where $H$ and $W$ represent the height and width of the input image, respectively. $\mathcal{P}_{max}$ is initialized to $H//2$. This design maintains spatial continuity in the row and column dimensions. It reduces the average distance between significant image tokens and instruction tokens, facilitating cross-attention among the image tokens and cross-attention between the image tokens and instruction tokens.
Subsequently, we propose gradually expanding the central receptive field to weaken the anchor tokens and enhance the model’s ability to perceive visual elements at varying levels of granularity. Specifically, we reduce $\mathcal{P}_{max}$ every $t$ layer, thereby controlling the perception granularity in terms of position encoding. When $\mathcal{P}_{max}$ is reduced to 1, the corresponding position encoding degenerates into All-one PE. In addition, we modify the causal mask $\mathcal{M}$ accordingly to ensure causal attention among tokens.
By introducing hierarchical position indexes, the Pyramid-descent Visual Position Encoding allows for a multi-granularity perception of visual elements. This method not only aligns better with human cognition but also improves the model’s overall performance in tasks that require both holistic and detailed perception of visual content.
@@ 4.2 Implementation Details
To demonstrate the generalizability of our proposed method across models with different parameter sizes, we conduct experiments using three model architectures with 3B, 7B, and 13B parameters. For 3B models, we follow TinyLLaVA Zhou et al. (2024) to use SigLIP Zhai et al. (2023) as the visual encoder and Phi-2 Li et al. (2023d) as the base LLM. For 7B and 13B models, we adopt pre-trained CLIP ViT-L/14 ($336^{2}$adford et al. (2021) as visual encoder and Vicuna v1.5 Zheng et al. (2023) as the base LLM. Following Liu et al. (2024a), we first pretrain the models on CC-558K dataset for 1 epoch with a learning rate of 1e-3, and finetune on the mix-665K dataset for 1 epoch, with a learning rate of 2e-5. All experiments are conducted on 8 NVIDIA A100 and 8 NVIDIA H20 GPUs.
@@ 5.2 Results of General Multimodal Benchmarks
As illustrated in Table 3, the PyPE series demonstrates exceptional performance on mainstream general multimodal benchmarks. In the MMStar benchmark, which is designed to assess genuine multimodal capabilities using visually indispensable samples, PyPE outperforms all baseline models. On MM-Vet, which evaluates the integration of core vision-language capabilities across 16 complex multimodal tasks, the 3B model of PyPE achieves an impressive score of 35.00, significantly surpassing the scores of 33.00 and 33.40 obtained by Raster-scan and Concentric PE, respectively. In the MMT-Bench evaluation, which assesses advanced reasoning and instruction-following across 32 core meta-tasks and 162 subtasks in multimodal understanding, PyPE markedly exceeds baseline performance, demonstrating its ability to apply expert knowledge and execute deliberate visual recognition, localization, reasoning, and planning. On MMBench, which evaluates fine-grained abilities across 20 dimensions, PyPE exhibits strong performance, matching or leading the state-of-the-art. Additionally, we test the methods on AI2D, a benchmark focusing on multiple-choice questions related to scientific diagrams containing text. The experimental results indicate that PyPE achieves state-of-the-art performance and demonstrates a strong comprehension of textual content within images.
5.3 Ablation Analysis
Effect of the Descent Interval
@@ 5.4 Qualitative Results on LLaVA-Bench
Figure 4 demonstrates a case study on how, given identical prompts and images, other baselines misperceive or inadequately process visual information, resulting in the generation of hallucinatory content. For instance, in the displayed example, the baseline methods exhibit object hallucinations, identifying nonexistent items such as "dining table", "hat", "scarf", and "boat". In contrast, the implementation of PyPE notably mitigates these hallucination issues while simultaneously maintaining the coherence and informativeness of the output text. This can be attributed to the multi-scale visual modeling capability afforded by the dynamic local receptive fields of PyPE, in conjunction with the stable global receptive fields. Furthermore, the visualization results of layer-wise attention indicate that our proposed method effectively alleviates the phenomenon of "aggregation pattern", thereby creating a synergistic effect with the former.
@@ 6 Conclusion
In this paper, we conduct an in-depth analysis of how visual position encoding affects visual perception in VLMs, particularly from the aspect of long-term decay and the "aggregation pattern". Our findings indicate that conventional visual position encoding methods are constrained by the "aggregation pattern" derived from Large Language Models (LLMs) and lacks multi-scale perceptual capabilities. To address these limitations, we introduce Pyramid-descent Visual Position Encoding (PyPE) a novel approach designed to enhance the perception of visual tokens within VLMs. Our extensive experiments across multiple benchmarks and VLM families demonstrate the efficacy of PyPE in addressing these challenges and ensuring a thorough understanding of visual content.
@@
Although PyPE demonstrates exceptional performance in enhancing the overall capabilities of Vision-language Models (VLMs), it is currently limited to single-frame images and has not yet been extended to video and other modalities. Future research will focus on how to effectively integrate the temporal dimension for unified position encoding and extending PyPE to a broader range of VLMs.

### v2

@@ 1 Introduction
Large Language Models (LLMs) Touvron et al. (2023); Brown et al. (2020) demonstrate significant universal capabilities that contribute to the pursuit of general artificial intelligence. However, language constitutes only one aspect of communication. Visual information plays a crucial role in augmenting and enhancing our understanding of the world. Consequently, there is a growing interest in the development of Vision-language Models (VLMs) Chen et al. (2024c); Peng et al. (2023); Wang et al. (2024); Bai et al. (2023) that can process and integrate visual modality. To effectively leverage the powerful contextual understanding capabilities of LLMs, VLMs project visual information to the same dimensionality as textual embeddings through specific projection layers Chen et al. (2023); Li et al. (2023b); Zhou et al. (2024), which are then directly embedded into the text sequence to form the input for the foundation LLMs, enabling cross-modal alignment and instruction-following learning using next-token prediction.
Despite their commendable progress, the typical processing of visual information does not align with the distribution patterns of visual elements. Since visual information is composed of fixed-sized patches obtained through raster scanning, patches located closer to the bottom right corner of the image are positioned nearer to the instruction tokens within the sequence. Due to the long-term decay from Rotary Position Embedding (RoPE) Su et al. (2024), visual tokens closer to the instruction tokens will be more likely to receive higher attention weights, and vice versa. This is counterintuitive, as the importance of visual information is not defined by the order of raster-scanning. Xing et al. (2024) observe a similar phenomenon by visualizing the attention information flow from instruction tokens to visual tokens in the first layer of the decoder. Consequently, they propose Concentric Causal Attention (CCA), which starts assigning the position indexes of images from the peripheral and ends in the center, to alleviate the long-term decay in RoPE and improve causal attention following 2D spatial locality of images. Although CCA is both intuitive and effective, its applicability is constrained by the assumption that all significant elements related to the instructions are situated at the center of the image. This assumption inherently results in a loss of detail, limiting its effectiveness in capturing comprehensive information.
@@ 2.1 Vision-language Model
Recent advancements in VLMs have demonstrated impressive performance in processing multi-format information Huang et al. (2023); Achiam et al. (2023). VLMs are typically built upon existing LLMs and incorporate visual information as input tokens by utilizing an additional vision encoder (e.g., CLIP) and a bridging connector (e.g., MLP). For instance, LLaVA Liu et al. (2024a) employs an MLP to project visual tokens and aligns the feature dimensions with word embeddings, while BLIP-2 Li et al. (2023b) utilizes a set of learnable query tokens to extract information in a query-based manner. Building upon these foundational works, MM1 McKinzie et al. (2025) has further investigated the significance of the number of visual tokens and image resolution, identifying them as the most critical factors, while finding that the type of connector has minimal impact. By effectively connecting visual and textual modalities, VLMs significantly enhance human-AI interaction and exhibit remarkable capabilities in understanding and generating multimodal content Chen et al. (2024b); Peng et al. (2023); Chen et al. (2023); Wang et al. (2024); Hu et al. (2024); Xie et al. (2024).
@@ 2.2 Position Encoding for Transformers
Since transformer-based models contain no recurrence Hochreiter (1997) and convolution Islam et al. (2020) structure, additional information about the relative or absolute position of the tokens in the input sequence is required. Therefore, the community has witnessed the development of various position encoding methods, e.g. sinusoidal Vaswani (2017), learnable Dosovitskiy (2020), relative He et al. (2020); Shaw et al. (2018), and conditional Chu et al. (2021) position encoding. Among these studies, RoPE Su et al. (2024) is introduced to encode absolute and relative positional information, showing superiority in LLMs Touvron et al. (2023); Achiam et al. (2023). The success of LLMs has led to the continued adoption of the effective RoPE scheme in VLMs for the unified encoding of positional information across sequences that incorporate multimodal features. However, it is important to note that visual information does not conform to the same sampling paradigm as language. The raster scanning is insufficient for modeling the spatial correlations among different patches. Consequently, numerous recent studies Chu et al. (2024); Xing et al. (2024); Lu et al. (2024) have sought to explore improved solutions that extend RoPE to visual tasks. In this paper, we investigate a novel multi-granularity position assignment strategy to enhance the VLM’s comprehension of visual information and improve the alignment between modalities.
@@ 3.1 Preliminaries
As indicated in Table 1, All-One PE performs weaker than the baselines in perception but keeps competitive in coarse-grained perception tasks on different sizes of models. This suggests that even when assigning the same positional weight to all image tokens, the VLM still possesses certain perception capabilities and performs better than raster-scan and concentric in coarse-grained situations. This is more pronounced on LLaVA-1.5-13B because larger models have stronger sequence modeling and feature capturing capabilities, which correspondingly bridge the gap in fine-grained abilities between All-One PE and other methods.
@@ 3.2 Pyramid-descent Visual Position Encoding
Similar challenges were also present in the early development of Vision Transformer (ViT) Dosovitskiy (2020). Due to the columnar structure of ViT, which uses coarse image patches as input, it is difficult to apply it directly to pixel-level dense predictions such as object detection and segmentation. This difficulty arises because its output feature map is single-scale and low-resolution. To address these issues, Wang et al. (2021) proposed the Pyramid Vision Transformer (PVT). They utilize fine-grained image patches as input to learn high-resolution representations and introduce a progressive shrinking pyramid to reduce the sequence length of the Transformer as the network deepens, significantly lowering the computational cost. Moreover, compared to CNNs, PVT consistently produces a global receptive field, ensuring a holistic perception of visual elements and benefiting its performance in detection and segmentation tasks.
@@
where $H$ and $W$ represent the height and width of the input image, respectively. $\mathcal{P}_{max}$ is initialized to $\lfloor H/2\rfloor$. This design maintains spatial continuity in the row and column dimensions. It reduces the average distance between significant image tokens and instruction tokens, facilitating cross-attention among the image tokens and cross-attention between the image tokens and instruction tokens.
Subsequently, we propose a gradual expansion of the central receptive field to diminish the influence of anchor tokens and enhance the model’s ability to perceive visual elements at varying levels of granularity. Specifically, we reduce $\mathcal{P}_{max}$ every $t$ layers, thereby controlling the granularity of perception through position encoding. When $\mathcal{P}_{max}$ is reduced to 1, the corresponding position encoding transforms into an All-One PE, which perceives more coarse-grained elements. To maintain causal attention, we adjust the attention mask $\mathcal{M}$ based on each assigned position matrix $\mathcal{P}$.
By introducing hierarchical position indices, PyPE facilitates multi-granularity perception of visual elements, allowing the model to dynamically adjust its focus to capture both broad contextual information and fine-grained details within visual data. This innovative approach not only aligns more closely with human cognitive processes but also enhances the model’s overall performance in tasks that require both holistic and detailed perception of visual content.
@@ 4.2 Implementation Details
To demonstrate the generalizability of our proposed method across models with different parameter sizes, we conduct experiments using three model architectures with 3B, 7B, and 13B parameters. For 3B models, we follow TinyLLaVA Zhou et al. (2024) to use SigLIP Zhai et al. (2023) as the visual encoder and Phi-2 Li et al. (2023d) as the base LLM. For 7B and 13B models, we adopt pre-trained CLIP ViT-L/14 ($336^{2}$adford et al. (2021) as visual encoder and Vicuna v1.5 Zheng et al. (2023) as the base LLM. Following Liu et al. (2024a), we pretrain the models on CC-558K dataset and finetune them on the mix-665K dataset. All experiments are conducted on 8 NVIDIA A100 and 8 NVIDIA H20 GPUs. See Appendix B for more training and implementation details.
@@ 5.2 Results of General Multimodal Benchmarks
As illustrated in Table 3, the PyPE series demonstrates exceptional performance on mainstream general multimodal benchmarks. In the MMStar benchmark, which is designed to assess genuine multimodal capabilities using visually indispensable samples, PyPE outperforms all baseline models. On MM-Vet, which evaluates the integration of core vision-language capabilities across 16 complex multimodal tasks, the 3B model of PyPE achieves an impressive score of 35.00, significantly surpassing the scores of 33.00 and 33.40 obtained by Raster-scan and Concentric PE, respectively. In the MMT-Bench evaluation, which assesses advanced reasoning and instruction-following across 32 core meta-tasks and 162 subtasks in multimodal understanding, PyPE markedly exceeds baseline performance, demonstrating its ability to apply expert knowledge and execute deliberate visual recognition, localization, reasoning, and planning. On MMBench, which evaluates fine-grained abilities across 20 dimensions, PyPE exhibits strong performance, matching or leading the state-of-the-art. Additionally, we test the methods on AI2D, a benchmark focusing on multiple-choice questions related to scientific diagrams containing text. The results indicate that PyPE achieves state-of-the-art performance and demonstrates a strong comprehension of textual content within images.
5.3 Analysis of the Descent Interval
@@ 5.4 Qualitative Results on LLaVA-Bench
Figure 4 demonstrates a case study on how, given identical prompts and images, other baselines misperceive or inadequately process visual information, resulting in the generation of hallucinatory content. For instance, in the displayed example, the baseline methods exhibit object hallucinations, identifying nonexistent items such as "dining table", "hat", "scarf", and "boat". In contrast, the implementation of PyPE notably mitigates these hallucination issues while simultaneously maintaining the coherence and informativeness of the output text. This can be attributed to the multi-scale visual modeling capability afforded by the dynamic local receptive fields of PyPE, in conjunction with the stable global receptive fields. Furthermore, the visualization results of layer-wise attention indicate that our proposed method effectively alleviates the phenomenon of "aggregation pattern", thereby creating a synergistic effect with the former. Refer to Appendix C for a more in-depth analysis of anchor tokens and Appendix E for more case studies.
@@ 6 Conclusion
In this work, we conduct an in-depth analysis of how visual position encoding affects visual perception in VLMs, particularly from the aspect of long-term decay and the "aggregation pattern". We find that conventional visual position encoding methods are constrained by the "aggregation pattern" derived from LLMs and lack multi-scale perceptual capabilities. To address these limitations, we introduce Pyramid-descent Visual Position Encoding (PyPE), a novel approach designed to enhance the perception of visual tokens within VLMs. Extensive experiments across multiple benchmarks and VLM families demonstrate the efficacy of PyPE in addressing these challenges and ensuring a thorough understanding of visual content.
@@
Although PyPE demonstrates exceptional performance in enhancing the overall capabilities of Vision-language Models (VLMs), it is currently limited to single-frame images and has not yet been extended to video and other modalities. Future research will focus on effectively integrating the temporal dimension for unified position encoding and extending PyPE to a broader range of VLMs.
@@ Appendix A Benchmarks
Figure 5: Visualization of anchor tokens in baselines and PyPE.
@@
     Hyperparameter      Pretrain      Finetune
     batch size      256      128
     lr      1e-3      2e-5
     lr schedule      cosine decay
     lr warmup ratio      0.03
     weight decay      0
     epoch      1
     optimizer      AdamW
     DeepSpeed stage      2      3
Table 5: Hyperparameters of TinyLLaVA-SigLIP-Phi-2 and LLaVA-1.5-7B/13B.
Method RefCOCO RefCOCO+ RefCOCOg
val test-A test-B val test-A test-B val test
TinyLLaVA-SigLIP-Phi-2
      w/ Raster-scan 31.85 15.77 44.29 31.28 18.65 40.83 56.86 56.98
      w/ Concentric 30.89 15.97 42.22 31.16 19.31 38.77 59.45 58.74
      w/ All-One 29.13 15.24 39.11 30.64 18.38 38.34 54.81 55.28
      w/ PyPE (Ours) 31.33 16.02 45.13 31.86 19.40 42.25 59.72 59.79
LLaVA-1.5-7B
      w/ Raster-scan 34.19 18.07 46.89 34.30 21.63 43.53 61.21 59.40
      w/ Concentric 32.23 16.51 42.49 32.66 20.00 40.41 59.72 58.47
      w/ All-One 32.99 16.46 41.26 33.28 20.73 39.83 63.07 61.90
      w/ PyPE (Ours) 35.16 16.51 48.04 34.17 21.22 41.46 64.62 64.13
LLaVA-1.5-13B
      w/ Raster-scan 36.86 19.29 50.01 36.12 22.37 43.59 63.66 60.96
      w/ Concentric 35.87 18.54 48.17 36.07 21.94 42.65 61.66 61.07
      w/ All-One 36.84 19.06 49.16 37.10 22.71 41.72 61.58 59.75
      w/ PyPE (Ours) 37.81 21.82 51.88 37.14 25.74 44.73 63.16 62.59
Table 6: Performance comparison on referring expression comprehension tasks. We use CIDEr Vedantam et al. (2015) to evaluate the quality of the descriptions. The highest results in each setting are indicated in bold, while the second-best results are underlined.
Appendix B Hyperparameters and More Implementation Details
We show the training hyperparameters for both first-stage vision-language alignment pretraining and the second-stage visual instruction tuning in Table 5. We use LMMs-Eval Zhang et al. (2024) to conduct experiments on VQA and general multimodal benchmarks.
Appendix C Visualization of Anchor Tokens
To further analyze the aggregating attention pattern, we visualize the attention score of each patch in the first 16 layers. As illustrated in Figure 5, both the All-One PE and the Concentric PE exhibit a relatively uniform distribution of attention in the initial two layers. However, a significant phenomenon of attention aggregation emerges in the subsequent layers, where non-anchor patches demonstrate a suppression of attention, particularly pronounced in Concentric PE. Though Raster-scan PE shows slight improvement, the attention in each layer tends to be preferentially allocated to patches that are closer to the instruction token, resulting in a discontinuous and fragmented attention pattern. This indicates a limitation of the Raster-scan PE in effectively modeling patches with similar semantics. In contrast, PyPE not only reduces the number of anchor tokens but also yields significantly lower attention scores for these tokens compared to the baselines, thereby facilitating the model’s exploration of image details more effectively. Furthermore, in each layer, the attention distribution of the PyPE is more continuous, highlighting the superiority of our proposed method in modeling semantically similar information.
Appendix D Performance on Referring Expression Comprehension
In the context of the visual localization task, we evaluate PyPE using the RefCOCO, RefCOCO+, and RefCOCOg datasets Kazemzadeh et al. (2014); Mao et al. (2016). The results, presented in Table 6, indicate that PyPE achieves top-tier performance among baselines. Its superior structural design enables PyPE to effectively perceive intricate details within images, resulting in significant improvements over baseline models. The performance of PyPE underscores its potential to advance the field of visual localization and its applicability in real-world scenarios that require precise visual understanding.
Appendix E More Case Studies
We provide more examples of visual description in Table 7. As illustrated in the table, our proposed PyPE exhibits a reduced incidence of generating visual hallucinations or misunderstandings. More importantly, compared to other baseline methods, PyPE demonstrates a finer granularity in perceiving visual elements, thereby uncovering additional information, such as "blueberries" in the first example and "My joke website (funny joke push to reveal punchline)" in the second example. To further analyze the model’s attention distribution across each decoder layer, we visualize the corresponding attention values for these examples. The results in Figure 67, and 8 indicate that while other baselines remain hindered by anchor tokens, PyPE consistently mitigates this issue, facilitating a more rational allocation of attention.
Visual input example, Visual Description 1
User Describe this photo in detail.
Raster-scan The image displays a refrigerator filled with various food items, including milk, juice, and fruits. There are multiple bottles of milk placed throughout the refrigerator, with some located on the top shelves and others on the bottom shelves. In addition to the milk, there are several bottles of juice, some of which are placed on the top shelves and others on the bottom shelves. A few carrots can be seen in the refrigerator, with some placed on the top shelves and others on the bottom shelves. The refrigerator is well-stocked with a variety of food items, making it a convenient place to store and access groceries.
All-One The image displays a refrigerator filled with various food items. Inside the refrigerator, there are several bottles, including one on the left side, two in the middle, and one on the right side. There are also multiple carrots, with some placed in the middle and others on the left side of the refrigerator. In addition to the carrots, there are several strawberries, with some located on the left side and others on the right side of the refrigerator. A bowl can be seen on the left side of the refrigerator, possibly containing more food items. The refrigerator is well-stocked with a variety of fruits and vegetables, making it a healthy and diverse selection.
Concentric The image displays a refrigerator filled with various food items, including a variety of fruits and vegetables. There are several carrots placed throughout the refrigerator, with some in the top left, middle, and bottom sections. A few strawberries can be seen in the bottom left corner of the refrigerator. In addition to the fresh produce, there are multiple bottles of milk and juice stored in the refrigerator. Some of these bottles are located in the top left, middle, and bottom sections of the refrigerator. The refrigerator is well-stocked with a diverse selection of food items, making it a convenient place to store and access groceries.
PyPE (ours) The image displays a refrigerator filled with various food items, including several containers of berries, milk, and other foods. There are multiple bottles of milk placed throughout the refrigerator, with some located on the top shelves and others on the bottom shelves. In addition to the milk, there are several containers of berries, such as strawberries and blueberries, spread across the refrigerator. The refrigerator is well-stocked with a diverse range of food items, making it a convenient place to store and access groceries.
Visual Description 2
User Describe this photo in detail.
Raster-scan The image features a handwritten sign with a joke written on it. The sign is placed on a table, and it appears to be a website or a piece of paper with a humorous message. The joke is written in a foreign language, which adds a unique touch to the sign. The sign is placed in the center of the image, drawing attention to its content.
All-One The image features a handwritten sign with a joke website written on it. The sign is placed on a white surface, possibly a piece of paper or a wall. The joke website is written in blue ink, and the text is accompanied by a question mark, indicating that the joke is a riddle or a puzzle. The sign is placed in the center of the image, drawing attention to the joke website.
Concentric The image features a handwritten sign with a joke on it, likely meant to be displayed on a website. The joke is written in a foreign language, and the sign is accompanied by a question mark, indicating that it is meant to be a humorous or amusing statement. The sign is placed on a white background, which contrasts with the dark blue writing.
PyPE (ours) The image features a handwritten sign with a joke on it, likely meant to be displayed on a website. The joke is written in blue ink, and it reads, "My joke website (funny joke push to reveal punchline)." The sign is placed on a white background, which contrasts with the blue ink used for the joke. The handwriting style suggests a casual and humorous tone, likely meant to entertain or amuse the viewers.
Table 7: More examples from LLaVA-Bench. The misunderstandings and hallucinations of visual contents are highlighted in red. The descriptions that are not mentioned in baselines but are accurately represented by PyPE are highlighted in green.
(e) Raster-scan
(f) Concentric
(g) All-One
(h) PyPE
Figure 6: Layer-wise attention visualization (left to right, up to down) of the example from Figure 4.
(a) Raster-scan
(b) Concentric
(c) All-One
(d) PyPE
Figure 7: Layer-wise attention visualization (left to right, up to down) of the first example from Table 7.
(a) Raster-scan
(b) Concentric
(c) All-One
(d) PyPE
Figure 8: Layer-wise attention visualization (left to right, up to down) of the second example from Table 7.